OpenStudy (anonymous):

Find the area of the region. Use a graphing utility to verify your result. y = 7 sin(x) + sin(7x)

OpenStudy (anonymous):

@ganeshie8 help?

OpenStudy (anonymous):

i got 14.286, but its wrong:(

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

this is in the wrong section, but that's fine \[\Large f ' (x) = 7\sin(x) + \sin(7x)\] \[\Large f(x) = \int (7\sin(x) + \sin(7x))dx\] \[\Large f(x) = \int (7\sin(x))dx + \int(\sin(7x))dx\] \[\Large f(x) = 7\int (\sin(x))dx + \int(\sin(7x))dx\] \[\Large f(x) = 7\int (\sin(x))dx + \frac{1}{7}\int(\sin(u))du \ ... \ u = 7x\] \[\Large f(x) = 7(-\cos(x)) + \frac{1}{7}(-\cos(7x))+C\] \[\Large f(x) = -7\cos(x) - \frac{1}{7}\cos(7x)+C\]

jimthompson5910 (jim_thompson5910):

wait, what endpoints does it give you?

OpenStudy (anonymous):

@jim_thompson5910 sorry i just saw it

OpenStudy (anonymous):

look

jimthompson5910 (jim_thompson5910):

which region are they referring to? what are the endpoints?

jimthompson5910 (jim_thompson5910):

ok from 0 to pi

jimthompson5910 (jim_thompson5910):

so you need to calculate f(pi) - f(0)

OpenStudy (anonymous):

thats exactly what i did and i got 14.286

OpenStudy (anonymous):

i also used wolfram to verify and it was the same answer

jimthompson5910 (jim_thompson5910):

write it as a fraction

jimthompson5910 (jim_thompson5910):

not the approximate answer

jimthompson5910 (jim_thompson5910):

what do you get?

OpenStudy (anonymous):

okay give me a second

OpenStudy (anonymous):

wolfram isn't giving me anything now. let me correct it

jimthompson5910 (jim_thompson5910):

do you see how I got \[\Large f(x) = -7\cos(x) - \frac{1}{7}\cos(7x)+C\]

OpenStudy (anonymous):

1/100

OpenStudy (anonymous):

100/7

OpenStudy (anonymous):

yes:)

OpenStudy (anonymous):

is 100/7 my answer then?

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (anonymous):

i see:) it was correct! thank you!

jimthompson5910 (jim_thompson5910):

you're welcome